# A Test-Time Learning Approach to Reparameterize the Geophysical Inverse Problem with a Convolutional Neural Network

### v1

@@ IV-C Enhanced Smoothing Effects from the Dropout Layer
Dropout is used to prevent the over-fitting issue in training large neural networks[20]. It surpasses conventional regularization methods such as Tikhonov regularization and max-norm regularization in supervised tasks in computer vision, speech recognition, etc[20]. When using dropout, we randomly zero out certain neurons during each training epoch; since different neurons are chosen to be zero-outed in each epoch, this is a stochastic regularization method. Intuitively, this stochastic zero-out scheme prevents the final output from relying too much on any particular neuron. Therefore, it improves the generalization of the neural network. If you view dropout as a model sampling method, in each epoch, it will sample a new CNN model with different zero-outed neurons than the model in the last epoch, so the output of the last epoch would be approximately a weighted average of different models[20]. See Fig. 13 and 14 for an ablation study on having/not having a dropout layer in Case 1 and 2, respectively. We choose a dropout rate of 0.1 and only add 1 dropout layer after the fourth hidden block of the DIP-Inv.
Figure 13: (a) is the true model for Case 1. b) is the inversion result from the proposed methods without a dropout layer, and c) is the inversion result from the proposed methods with a dropout layer.
Figure 14: (a) is the true model for Case 2.1. b) is the inversion result from the proposed methods without a dropout layer, and c) is the inversion result from the proposed methods with a dropout layer.
The dropout layer is useful for improving the recovery of the top layer of the subsurface for Case 1 as shown in Fig. 13 and recovery of the cylindrical target for Case 2 as shown in Fig. 14. A practical guide for applying the proposed method would be to perform the proposed method without the dropout layer in the first attempt, and then perform the proposed method with the dropout if the first result is not satisfying.
@@ V Future Work
The idea of reparameterizing the mesh space can be adapted to other Tikhonov-style geophysical inversion problems (e.g. potential fields and electromagnetics, where SimPEG has other simulations that could be swapped in.). This method can also be combined with additional explicit regularization terms such as using a non-uniform reference model based on the petrophysical knowledge for the survey domain. Currently, the DIP-Inv method requires more than 1000 iterations, so it is much slower than a conventional inversion. Exploring more efficient optimizers is a direction of future research. Moreover, to illustrate the impacts of the implicit regularization on models with known structures, we only show the simple synthetic cases. Future work could also include examining the results of field data. Similarly, we chose to work with a simple CNN architecture, but exploring more complex architectures would be an avenue for further research.
@@ VI Conclusion
Motivated by the work of utilizing the implicit regularization in the inverse problems in other fields, we explore the utility of the implicit regularization inherently included in the CNN structure for the Tikhonov-style geophysical inversion. In this study, we focus on the DC resistivity with synthetic examples and compare the DIP-Inv results with the conventional Tikhonov-regularized inversion results using the sparse norms. Compared to the conventional method, the DIP-Inv method improves the recovery of the dip of a target and the recovery of the compact targets in close proximity. We also show that the bi-linear upsampling operator is a key component in DIP-Inv and adding dropout can also have advantages in some cases. Those results illustrate that the implicit regularization from the CNN structure can be beneficial for the Tikhonov-style geophysical inversion. As a test-time learning method, DIP-Inv can be adapted to other geophysical inversion problems or combined with explicit regularization terms.
@@ Acknowledgments
We thank Jorge Lopez-Alvis for his comments and suggestions on paper writing. We would also like to thank Doug Oldenburg, Devin Cowan, Joseph Capriotti, John M. Weis, Johnathan C. Kuttai, Jingrong Lin and Santiago Soler for their invaluable guidance and advice on the use of SimPEG.
Details of the DIP-Inv
As we mentioned in the paper, the dimension of $y_{1}$ and other blocks should change to adapt the mesh design. For the two synthetic cases, where they share the same mesh design, we list the details of the hidden blocks in the Table below.
TABLE I: Details of the architectures implemented in the two synthetic cases
hb1 FC (8, 174)
LeakyReLU (negative slope 0.2)
reshape (1,1,6,29)
hb2 Bi-linear upsampling ($k=2$)
Convolutional ((1,64),
kernel ($3\times 3$),
stride (1),
padding (0))
LeakyReLU (negative slope 0.2)
hb3 Bi-linear upsampling ($k=2$)
Convolutional ((64, 32),
kernel ($3\times 3$),
stride (1),
padding (0))
LeakyReLU (negative slope 0.2)
hb4 Bi-linear upsampling ($k=2$)
Convolutional ((32, 8),
kernel ($3\times 3$),
stride (1),
padding (0))
LeakyReLU (negative slope 0.2)
hb5 Convolutional ((8, 1),
kernel ($3\times 3$),
stride (1),
padding (0))
Sigmoid
output cropped, ×8$\times$-8× - 8
Appendix A Biography Section
[Uncaptioned image] Anran Xu received the Honours Bachelor of Science degree in Mathematics &\&& Its Applications Specialist (Physical Science) and Physics Major from the University of Toronto, Toronto, ON Canada, in 2022. She is currently pursuing a Master of Science in Geophysics with the University of British Columbia, Vancouver, BC Canada. Her research interests include inverse problems and machine learning applications.
[Uncaptioned image] Lindsey J. Heagy is an Assistant Professor in the Department of Earth, Ocean and Atmospheric Sciences and Director of the Geophysical Inversion Facility at UBC. She completed her BSc in geophysics at the University of Alberta in 2012 and her PhD at UBC in 2018. Prior to her current position, she was a Postdoctoral researcher in the Statistics Department at UC Berkeley. Her research combines computational methods in numerical simulations, inversions, and machine learning to characterize the geophysical subsurface.

### v2

@@ IV-C Enhanced Smoothing Effects from the Dropout Layer
Dropout is used to prevent over-fitting in training large neural networks [42]. It surpasses conventional regularization methods such as Tikhonov regularization and max-norm regularization in supervised tasks in computer vision, speech recognition, etc [42]. When using dropout, we randomly zero out certain neurons during each training iteration. Since different neurons are chosen to be zero-outed in each iteration, this is a stochastic regularization method. Intuitively, this stochastic zero-out scheme prevents the final output from relying too much on any particular neuron. Therefore, it improves the generalization of the neural network. Dropout can be viewed as a model sampling method as well. In each iteration, it will sample a new CNN model with different zero-outed neurons compared to the model in the last iteration, so the output of the last iteration would be approximately a weighted average of different models [42]. See Fig. 15 and 16 for an ablation study on having/not having a dropout layer in Case 1 and 2, respectively. We choose a dropout rate of 0.1 and only add 1 dropout layer after the fourth hidden block of the DIP-Inv.
Figure 15: (a) is the true model. (b) is the inversion result from the DIP-Inv methods without any dropout layer, and (c) is the inversion result from the DIP-Inv methods with a dropout layer.
Figure 16: (a) is the true model. (b) is the inversion result from the DIP-Inv methods without any dropout layer, and (c) is the inversion result from the DIP-Inv methods with a dropout layer.
The dropout layer is useful for improving the recovery of the top layer of the subsurface for Case 1 as shown in Fig. 15 and recovery of the cylindrical target for Case 2 as shown in Fig. 16. A practical guide for applying the DIP-Inv would be to perform it without any dropout layer in the first attempt, and then perform it with the dropout layer if the first result is not satisfying.
@@ V Future Work
The above analyses are mainly based on experimental results, a more rigorous mathematical explanation is a natural direction of the future work. The mathematical explanation for the implicit bias of neural networks is still an open question. For example, Frei et al. [43] proved that a 2-layer Leaky ReLU MLP is a $l_{2}$-max-margin solution under certain conditions. Rahaman et al. [44] found that ReLU NN is biased toward learning lower-frequency variation. Some scholars have studied the implicit bias from the gradient-based optimization methods to answer the generalization puzzle in NN [45]. Specifically, for convolutional neural networks, Tachella et al. [46] stated that Adam is key to the success of utilizing the DIP method in image denoising. These explanations are limited by the assumptions they made to prove the relevant theorems and do not perfectly interpret the implicit regularization effects that we have observed in DIP-Inv. However, these works shed light on explaining the good performance of over-parameterized neural networks.
The idea of reparameterizing the mesh space can be adapted to other Tikhonov-style geophysical inversion problems (e.g. potential fields and electromagnetics, where SimPEG has other simulations that could be swapped in). This method can also be combined with additional explicit regularization terms such as using a non-uniform reference model based on the petrophysical knowledge for the survey domain. Currently, the DIP-Inv method requires more than 1000 iterations, and it is much slower than a conventional inversion, so exploring more efficient optimization methods is also a direction of future research. Moreover, to illustrate the impacts of implicit regularization on models with known structures, we only show the synthetic cases. Future work could also include examining the results of field data. Similarly, we chose to work with a simple CNN architecture, but exploring more complex architectures would be an avenue for further research.
@@ VI Conclusion
Motivated by the works of utilizing the implicit regularization inherently included in the CNN structure in the inverse problems in other fields, we explore the utility of this implicit regularization in the Tikhonov-style geophysical inversions. In this study, we focus on the DC resistivity inversion with synthetic examples and compare the DIP-Inv results with the conventional Tikhonov-regularized inversion results using the sparse norms. Compared to the conventional method, the DIP-Inv method improves the recovery of the dip of a target and the recovery of the compact targets in close proximity. We also show that the bi-linear upsampling operator is a key component in DIP-Inv and adding dropout can be beneficial in some cases. Those results illustrate that the implicit regularization from the CNN structure can be useful for the Tikhonov-style geophysical inversion. As a test-time learning method, DIP-Inv can be adapted to other geophysical inversion problems or combined with explicit regularization terms without collecting a training dataset.
@@ Acknowledgments
This research was supported by funding from the NSERC Discovery Grants Program and the Mitacs Accelerate Program. We thank Jorge Lopez-Alvis for his comments and suggestions on paper writing. We would also like to thank Doug Oldenburg, Devin Cowan, Joseph Capriotti, John M. Weis, Johnathan C. Kuttai, Jingrong Lin and Santiago Soler for their invaluable guidance and advice on the usage of SimPEG.